OpenStudy (anonymous):

solve each system of equations -6v-6y=66 2v+4y=-32

OpenStudy (ryan123345):

What do u want to use substitution or elimination?

OpenStudy (anonymous):

We can solve such system of equations by making any of the variables equal and then subtracting it from either of the equations

OpenStudy (anonymous):

elimination works better for me I think

OpenStudy (anonymous):

like multiply second equation with 3 3 x (2v+4y)=-32 x 3 6v + 12y = -96

OpenStudy (anonymous):

now add both the equations 6v + 12y = -96 -6v - 6y= 66 0 + 6y = - 36 y = -6

OpenStudy (anonymous):

You can put this value of Y = -6 in any of the equation to get the value of v HOPE you can solve from here

OpenStudy (anonymous):

just put y = -6 in of the equation..suppose in equation 1 -6v-6y=66 -6v + 36 = 66 -6v = 30 v = - 5